### v2

# New Online Communities: Graph Deep Learning on Anonymous Voting Networks to Identify Sybils in Polycentric Governance

Abstract
This research examines the polycentric governance of digital assets in blockchain-based Decentralized Autonomous Organizations (DAOs). It offers a theoretical framework and addresses a critical challenge facing decentralized governance by developing a method to identify sybils, or spurious identities. The method uses graph deep learning techniques to identify sybil activity in a DAO governance dataset (snapshot.org). Specifically, a Graph Convolutional Neural Network (GCNN) learned voting behaviours and a fast k-means vector clustering algorithm (FAISS) used the high dimensional embeddings to identify similar nodes in a graph. The results reveal that deep learning can effectively identify sybils, reducing the voting graph by 2-5%. This research underscores the importance of sybil resistance in DAOs and offers a novel perspective on decentralized governance, informing future policy, regulation, and governance practices.
1   Introduction
Decentralized Autonomous Organizations (DAOs) represent a critical intersection of social science and emerging technologies. Originating from blockchain architectures, DAOs have transcended their initial role in cryptocurrency governance to become complex social systems that defy traditional classification. In an attempt to elucidate their underlying structure and functions, this research offers a theoretical framework for DAOs and presents an inductive method to identify sybils, thereby tackling a substantial issue facing the decentralized governance of digital assets.
In this research, a theoretical exploration of endogenous and exogenous models drawn from research on Online Communities, coupled with empirical data, shapes the foundation of the neoinstitutional theory offered here. In a nod to Elinor Ostrom’s work, this theory posits that DAOs can effectively govern Digital Common Pool Resources (DCPRs) through polycentric governance, despite their idiosyncratic challenges, if they effectively manage sybils. Sybils are fraudulent identities used to manipulate voting within digital decentralized governance. While traditional governance frameworks readily address the ‘one person, one vote’ principle, the anonymity provided by blockchain technologies complicates the influence of sybils in decentralized governance.
To address this governance challenge, we used graph deep learning techniques to identify sybil activity on a popular DAO governance forum, snapshot.org. Specifically, our method involves a Graph Convolutional Neural Net (GCNN) with multiple engineered feature layers that learned vector node embeddings from a voting graph to construct a “similarity” subgraph, where similar nodes are identified using a fast k-means clustering algorithm (Facebook AI Similarity Search, FAISS) and relabelled. The similarity graph is then reduced by combining predicted sybils. While this method is highly inductive and sensitive to feature development and learning hyperparameters, experimental results suggest that deep learning techniques can meaningfully detect sybils and the resulting subgraph of ‘true’ identities is approximately 2-5% reduced. Social scientists, policy makers, system designers, and security professionals should consider these results when studying, designing, and implementing online communities and ensure that vulnerable governance practices are made resilient to intentional and unintentional sybil attacks.
2   Deciphering DAOs: Advancing Research in Digital Collective Action
The study of Decentralized Autonomous Organizations (DAOs) is marked by a burgeoning empirical landscape, rich with detailed quantitative analyses. This section focuses on the research challenges that confront DAO Studies, underscored by the nascent nature of the field, the lack of established quantitative baselines for comparison, and the complexities introduced by anonymous networks and sybil entities.
The fluidity and lack of standardization in DAOs challenge conventional theories and frameworks of organizations, institutions, and online communities. Far removed from the hierarchical corporate models of the last century, DAOs are more like distributed, loosely coupled networks often seen in non-hierarchical collectives like social movements or even terrorist cells. A focus on their decentralized and autonomous nature, however, can be misleading; in practice, DAOs frequently fail to achieve either decentralization or autonomy in an unqualified sense.
Despite their inherent instability, variability, and quasi-anonymous infrastructure, DAOs are not beyond the scope of systematic study. In fact, their wide-ranging goals, from digital ‘vibes’ to addressing complex real-world goals, highlights the need for scholarly attention. We have observed that increasingly new platforms are imagined, designed, built, and maintained by communities of loosely connected participants who have an economic stake in the outcome, rather than by “core” teams of developers with commit access, as with Bitcoin, Ethereum, and earlier centralized Web2/FLOSS platforms like Wikipedia or Linux. This makes DAOs both similar to yet distinct from existing Online Communities, a subfield of study for Information Systems.
The key unifying factor among all DAOs is their use of, or responsibility for, a decentralized ledger, or “blockchain.” Central to each DAO is a shared digital asset on a blockchain, often referred to as a “token.” Recognizing the pivotal role of “token economics” in these novel, community-led platforms, we find a compelling area of research: the exploration of how these digitally augmented human collectives manage these shared assets. This exploration draws our attention towards the concept of a “Common Pool Resource” [35] and my suggested adaptation, the “Digital Common Pool Resource” (DCPR). The study of DCPRs, as a result, could offer valuable insights into the nature and operation of DAOs.
The importance of DAO research also goes beyond academic interest. It offers an opportunity to challenge and extend our understanding of collective action, digital resource management, and decentralized governance. Further, studying DAOs offers insights into ongoing, cutting-edge experiments attempting to create robust, long-enduring institutional forms beyond the control (and regulation) of nation states. Moreover, these efforts buttress new and existing cryptocurrencies and frustrate criminal investigations, which routinely seek cryptocurrency traces when investigating crime. More optimistically and progressively, through detailed data analysis we can better comprehend the intricacies of DAOs and tap into their potential to reshape our digital environments, perhaps even as a form of “plural public” [28].
Before describing how the emergence of Digital Common Pool Resources (DCPRs) extends the existing neoinstitutional, resource-management perspective, we offer a brief literature review of the emerging subfield of “DAO Studies” by explicating existing empirical measurements and detailing the conceptual challenges these studies face. Then we briefly connect this perspective to Online Communities literature in Information Systems, in effect linking previous studies on “Web 2.0” to a new phenomenon in “Web3.” Informed by both the dynamics of online communities and commons resource management practices, we gesture towards a theory of digital asset management in a true digital commons. We conclude with an experimental analysis of a large DAO dataset (snapshot.org) by developing a graph deep learning method for detecting sybils, addressing a key existing challenge to successful decentralized governance.
3   Emergence and Evolution of DAO Studies
In the early 2020s, a nascent, specialized field of “DAO Studies” emerged (see [50]). This examination of what we now call “Web3” offers a new perspective on the understanding and analysis of digital collectives. While DAOs have existed formally since the emergence of the Ethereum platform in early 2015, their first year of operation was rarely studied and few, even in the industry, appreciated their eventual significance. The trajectory of this field took a definitive turn following a seminal event: the 2016 “The DAO” attack.
This event, which exposed vulnerabilities of DAOs but also spurred interest in, and the development of, decentralized governance technologies, brought about a significant shift in the discourse surrounding these structures. The 2016 The DAO attack (see [14]) serves as a key event, illuminating emerging issues surrounding the governance and legitimacy of DAOs and contributing to the theoretical foundations of a rapidly emerging DAO Studies.
Following a critical examination of this event, the field of DAO Studies started to grapple with endogenous, actor-produced theories in ways that highlighted key issues like decentralization, legitimacy, and the transformation of trust. Rooted in the experiences and perspectives of key actors in DAOs, these endogenous theories attempted to capture the often utopian vision, or social “imaginary,” that undergirds the diverse communities. The emergence of such a perspective, outlined by Dylan-Ennis [16] and others, signifies a move towards a more nuanced understanding of the aspirations and ideologies that shape DAOs but remains locked to endogenous, actor-produced theories.
As the discourse within DAO Studies evolves, we are starting to see a shift from insular, participant-derived narratives towards the integration of established socio-economic theories that offer a broader context. A pivotal theoretical framework used to analyze these phenomena has been Hirschman’s paradigm of “Exit, Voice, and Loyalty” [27]. In this model, ‘Exit’ refers to the option of leaving an organization or community when its quality deteriorates or when disagreements arise, while ‘Voice’ involves expressing dissatisfaction or working towards change from within the organization. ‘Loyalty,’ on the other hand, reflects a commitment to the organization that might inhibit ‘Exit’ and promote ‘Voice.’
In the context of DAOs, ‘Exit’ corresponds to members’ ability to disassociate from the organization, typically by selling their tokens, in response to dissatisfaction or perceived decrease in value or utility. In comparison, ‘Voice’ manifests in the community’s decision-making processes, where token holders can vote on proposals or express their opinions in fora, thereby actively participating in shaping the organization’s trajectory. The traditional notion of ‘Loyalty,’ however, undergoes a transformation in the realm of DAOs. With the advent of ‘Exit to Community’ (E2C) – a concept coined by Schneider and Mannan [44] – ‘Loyalty’ can morph into a mechanism for community ownership. E2C emphasizes that a loyal user or member, rather than simply exiting when dissatisfied, may have the option to help transform the organization into a more community-oriented or user-owned entity, thus introducing a novel form of ‘Loyalty’ that redefines the relationship between community members and the DAO.
Similarly, many scholars have addressed how trust is transformed by DAOs. Much literature on the role of trust in blockchain technology presupposes that its function is to lower transaction costs within organizations, a sentiment summarized by Williamson [52] with his famous theory of Transaction Cost Economics (TCE). Halaburda et al.’s [24] research extends this argument by exploring how blockchain, as a trust-enhancing technology, can alter the dynamics of transaction governance in organizations, including DAOs. This work explores the intricacies of how blockchain’s inherent features, such as decentralization, transparency, and immutability, contribute to reducing uncertainties and transaction costs that are traditionally managed through hierarchical organizational structures. Halaburda’s insights are particularly relevant in the context of DAOs, which are thought to operate on principles of decentralized governance and collective decision-making. By reducing the need for intermediaries and centralized control, blockchain technology enables DAOs to operate with a higher degree of trust among participants, fostering a collaborative environment that is less susceptible to opportunistic behaviors, a key concern in TCE.
Moreover, Halaburda’s perspective aligns with Hardin’s theory of trust, emphasizing how trust in DAOs is not merely a function of self-interest or moral commitment but also heavily influenced by the technological framework that underpins these organizations. This viewpoint is further developed by Lemieux [29], who reconstructs Hardin’s theory [26] to incorporate the role of blockchain technology in shaping trust dynamics. A different model of trust emerges from Hardin’s theory of trust, where trust is “grounded in the truster’s assessment of the intentions of the trusted with respect to some action,” which is, typically, based in self-interest, moral commitment, and other idiosyncratic factors. These motivations are clearly present in the organizational dynamics of DAOs, however without alteration Hardin’s theory fails to address the increasing role of technology. Others, such as Wright and De Filippi [53], have focused on the ways technology transforms trust, drawing a parallel to the substitutability, or lack therefore, of law by code.
A central theme of these theories, and indeed within the broader field of DAO Studies, is the importance of decentralization. Motivations driving DAOs towards decentralization are everywhere. They encompass the quest for enhanced security and robustness; oftentimes there is an ideological commitment to political decentralization; and in the pockets of progressive Web3, a desire to foster a global, diverse community. As well, the discourse on decentralization naturally segues into an exploration of legitimacy and the transformation of trust, as just discussed.
However, some scholars question whether decentralization is really the right term for the communities’ goals (for example, JP Vergne [49] prefers “dispersion of authority”). Others have approached the question of decentralization by measuring variables believed to be representative of an individual’s control over change. Indeed, decentralization has been a key measurement in the early quantitative and empirical literature attempting to understand DAO governance dynamics.
3.1   Quantitative and Empirical DAO Research
The empirical landscape of DAOs has been significantly expanded by several key studies, each contributing to a deeper understanding of DAO governance and operational dynamics.
Liu et al.’s [30] research provides an analysis of 50 DAOs, selected based on treasury size, number of token holders, and proposal frequency. The study’s multi-method approach, examining proposal texts from five DAOs and studying voting power, reveals a concerning voter turnout of merely 1.77%, challenging the expected participatory norms of DAOs. This quantitative detail underpins the assertion of centralized control within DAOs, further quantified by the disproportionate voting power held by top players—35% by the top player alone, and even more concentrated when considering the collective influence of the top three players, who control 63% of the voting power.
Wang et al.’s [50] empirical study spans a dataset of 581 DAOs with 16,246 governance proposals. Their methodology, utilizing k-means clustering and Principal Component Analysis (PCA) on proposal titles from 2020 to 2022, brings to light the difficulty of achieving decentralization. This study’s approach, which used data scraping instead of API access, reasserts Snapshot’s claim of a 95% coverage of DAOs, although it raises critical methodological questions, such as the accurate counting of sybils within member counts.
Sun et al. [47] analyzed MakerDAO’s voting from August 5, 2019, to October 22, 2021. Their empirical analysis introduced three measurements for centralized governance: voting participation, centralized voting power via the Gini coefficient, and the largest voter’s ‘voting power.’ Their results indicate a pronounced impact of centralized governance on MakerDAO, affecting financial stability, transaction volumes, network growth, and public perception as also reflected in Twitter sentiment analyses.
Sharma et al. [45] conducted a study involving ten DAOs, blending qualitative insights from interviews with quantitative on-chain data analysis. This multi-method approach found significant centralization of voting power; in particular, CompoundDAO, where out of 2,482 unique addresses, 32 ‘whale’ addresses hold between 100,000 to 10 million COMP tokens each, significantly swaying proposal outcomes.
Rikken et al. [42] analysed 4,592 DAOs, distinguishing 3,881 with measurable activity. Their categorization is based on activity level, token holders, and deployment types, supported by a survival analysis through Receiver Operating Characteristic (ROC) curves. They find that DAOs with at least 20 token holders are likely to have a higher survival rate. The study ventures beyond mere survival indicators, engaging with emerging trends such as off-chain voting, and provocatively suggested that certain DAOs might be more accurately described as Decentralized Organizations due to these limitations.
Goldberg and Schär’s [23] study of Decentraland’s DAO governance analyzes 1,414 governance proposals, involving 45,333 votes from 4,345 addresses, to elucidate the ‘whale effect’ and its influence on governance centralization. Their research revealed critical disparities in voting power and the strategic timing of votes, which impact the democratic integrity of the governance process.
Two further studies on MakerDAO reveal its importance to the decentralized governance of digital assets. Sun, Stasinakis, and Sermpinis [46] conducted a detailed investigation into voter coalitions within MakerDAO, utilizing a k-means clustering algorithm on 809 governance polls, uncovering the impact of voter coalitions on protocol stability and growth. Zhao et al. [56] bring an Organizational Management perspective to task management in MakerDAO, with a binary variable analysis of strategic versus operational voting among 378 proposals. Echoing others, their findings present challenges to decentralization due to the dominance of powerful stakeholders, a concern echoed across DAO studies.
Faqir-Rhazoui, Gallardo, and Hassan [18] present a comparative analysis of decentralized governance adoption across three main DAO platforms—Aragon, DAOstack, and DAOhaus—using data from 72,320 users across 2,353 DAO communities to evaluate growth, activity, voting systems, and funds management.
Xu et al. [55] and Fan et al. [17] investigate different aspects of DAO functionality—automated decision-making and role identification, respectively. Xu et al. developed a governance bot using reinforcement learning for automation or decision support of DAO governance. Fan et al. analyzed 20,148 accounts on the popular DeFi platform Paraswap, categorizing transactional behaviours to identify participant roles and community dynamics, and highlighted the influence of speculative activities.
In an effort to understand and categorize the diversity of forms DAOs take, Wright [53] proposed methodologies for measuring DAO autonomy by drawing lessons from other autonomous systems. Axelsen, Jensen, and Ross [3] examined organizational typologies in DAOs to determine when a DAO can be deemed decentralized. Ziegler and Welpe [58] developed a taxonomy for DAOs by exploring 35 DAOs and used agglomerative clustering to categorize them into five clusters. Finally, Braun, Haeusle, and Karpischek [8] attempt to model DAOs using a Transaction Cost Economics (TCE) perspective and consider the impact of revenue sharing, incentives for staking (investing), decentralization, and a variety of voting schemes. Like others who considered the role of decentralization, they model decentralization in terms of an Herfindahl-Hirschman Index (HHI), which is formally similar to GINI calculations [45][47] and top-player [30] and whale analyses [23].
Herfindahl-Hirschman: HHI =
$\sum_{i=1}^{N}s_{i}^{2}$
Where $s_{i}$ is the market share of firm $i$ and $N$ is the total number of firms.
Gini: G =
$\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|x_{i}-x_{j}|}{2N^{2}\bar{x}}$
Where $x_{i}$ and $x_{j}$ are the values of the income or wealth of individuals $i$ and $j$, $N$ is the number of individuals, and $\bar{x}$ is the mean income or wealth.
$n$-top Players:
$C_{n}=\sum_{i=1}^{n}s_{i}$
Where $C_{n}$ is the concentration of the top $n$ players, $s_{i}$ is the market share of the $i$-th firm, and $n$ is the number of top firms being considered.
3.2   Existing Research Challenges for DAO Studies
DAO research is an emergent field, characterized by the absence of standardized baselines that facilitate comparative analysis. This lack of common benchmarks hinders the ability to assess the efficacy and impact of DAO governance models against established norms. Furthermore, the field is yet to develop a unified institutional theory that can frame a common set of issues and concerns for researchers to address.
The theoretical underpinnings of DAO studies are relatively undeveloped, with the initial body of literature often drawing from a ‘native’ theory of technological utopianism that is self-referential and uncritical. This self-referential use of theory may have lead to an echo chamber effect that focused attention on decentralization and autonomy at the expense of richer social phenomena.
A significant challenge in DAO research is the anonymized nature of blockchain networks. Sybils, or duplicate identities controlled by the same entity, compound this problem, making it difficult to distinguish genuine community engagement from manipulated governance activities. The prevalence of sybils within DAO structures can distort empirical analyses and mislead interpretations of data, resulting in flawed conclusions about the state of decentralization and participant behaviour.
These challenges necessitate the development of sophisticated methodologies and research tooling tailored to the unique characteristics of DAOs. Without addressing these methodological and theoretical gaps, the field may struggle to produce reliable and actionable insights that can inform the evolution of DAOs and their governance frameworks. Therefore, the future of DAO research lies in overcoming these hurdles to pave the way for more robust, comprehensive, and accurate studies within this dynamic and rapidly evolving domain.
4   Online Communities
Mindel et al.’s [33] study provides valuable insights into the dynamics of online communities by applying the polycentric governance principles derived from Elinor Ostrom’s work to Web2 online communities. The study extends the original concept of polycentric governance to digital platforms, shedding light on how decentralized systems can benefit from community-driven governance for sustainable development and resource management. Here, we want to further explore polycentric governance and align it with the challenges faced by DAOs. DAOs, as decentralized platforms, require effective management of resources and engagement without central authority. Mindel et al.’s exploration into the sustainability of polycentric information commons offers a framework that could be instrumental in addressing the new challenges facing the management of digital asset commons in Web3.
In decentralized online information systems, the principles of polycentric governance help counterbalance sustainability threats posed by unrestricted participation (which may lead to Hardin’s famous “Tragedy of the Commons” [25] if left unchecked). These principles, manifested in system rules and infrastructure features, are crucial to understand how some systems decline while others flourish despite collective action threats inherent in their openness. Mindel et al.’s relevance to DAOs lies in their detailed analysis of various factors that impact the sustainability of online platforms. These include the dynamics of provision, appropriation, revitalization, and equitability, as well as collective action threats like free-riding, congestion, pollution, violation, and rebellion. By understanding these dynamics, DAOs can implement governance structures that effectively manage community contributions, maintain equitable participation, and mitigate threats to their sustainability.
4.1   Motivation and Governance in the Transition from Web 2.0 to Web3
Engagement in online communities has evolved from the simple pleasures of Web 2.0 to the complex incentive structures of Web3. Linus’s law, as proposed by Raymond [41], captured the essence of early online collaboration, suggesting that peer-driven development not only yields effective results but is also inherently enjoyable. Expanding on this, Dejean [12] emphasized that the allure for participants often lies in the intrinsic rewards of community engagement and recognition, over monetary gains. This perspective is increasingly relevant in the decentralized governance framework of DAOs, where traditional corporate rewards are replaced by token incentives and collective decision-making processes, as analyzed by De Filippi and Hassan [21].
The dichotomy of motivation in online communities is well-established, with Faraj et al. [19] showing that both intrinsic and extrinsic factors are critical in fostering engagement. In DAOs, smart contracts quantify and enforce commitment, potentially aligning personal motivations with communal goals. Gallus [22] indicates that symbolic rewards in communities can enhance commitment. Yet, the unique financial dynamics within DAOs complicate this simple dichotomy and underscore the need for empirical scrutiny to understand the interplay between these motivators.
Knowledge sharing and collaborative production are the lifeblood of open-source ecosystems, epitomized by Linux and Wikipedia, which stand as testaments to the power of collective intelligence. In the emergent landscape of Web3, DAOs represent a transformative expansion of these principles, infusing them with advanced decentralization and economic mechanisms. Furthermore, these structures enable novel revenue-sharing and ownership models that equitably distribute the fruits of collective efforts, challenging traditional paradigms of ownership and profit allocation.
This evolution in the roles of contributors within DAOs is not merely a change in operations but signifies a shift in the underlying ethos of organizational structures. It heralds an era where the value generated by a community is returned to it, fostering a more engaged, motivated, and equity-driven environment. Thus, DAOs do not just adopt the foundational aspects of open-source projects; they adapt and amplify them, introducing a new model of collective action where governance and gains are intrinsically linked to participation.
The renewed focus on social capital in online communities and DAOs becomes a cornerstone for innovation and trust. Dahlander [11] discusses the impact of multiple affiliations on innovation, which is relevant in DAOs where decentralized control of digital assets creates new collaboration patterns. This challenges the previously held primacy of network effects, a shift underscored by Faraj et al. [19].
Behavioral dynamics and social roles in online communities have been explored by researchers such as Benamar [5], but DAOs introduce a shift in these structures. Smart contracts facilitate the reimagining of roles, leading to novel governance structures within DAOs that depart from the rigidity of traditional online communities, allowing more adaptive leadership and participation models, as explored in Web 2.0 by Carillo [9].
Lastly, and key to our investigation, the shift to decentralized communities presents new governance challenges. Issues such as sociability and fringe behavior, often overlooked in research, are integral to DAOs and necessitate innovative governance frameworks, a concept echoed by Faraj, Srinivas, and Wasko [32]. DAOs serve as a testing ground for new approaches to leadership, community engagement, and the valuation of communal resources, representing a significant departure from the governance models of traditional online communities.
5   Towards Polycentric Governance
Elinor Ostrom, in her seminal work Governing the Commons [35], introduced a neoinstitutional approach that challenged the notion that common pool resources (CPRs) are inevitably doomed to depletion, as suggested by the “Tragedy of the Commons” [25]. The “Tragedy of the Commons” refers to the idea that individuals, acting in their self-interest, will deplete shared resources, leading to a loss for the entire community. This concept, first put forth by Garrett Hardin, has been the subject of extensive scholarly debate and analysis. Hardin himself later clarified that his focus was on “unmanaged commons” [26], and scholars such as Feeny et al. [20] have explored various common-pool resource dilemmas, emphasizing the diversity of outcomes depending on governance structures.
Nonetheless, Ostrom demonstrated through empirical evidence how community-based management, rather than top-down control, could effectively govern shared resources. Her approach went beyond Hardin’s conceptualization, as she identified critical factors that enable communities to sustainably manage resources, contrary to Hardin’s more deterministic view. She challenged Hardin’s view through extensive case studies, such as fishing in Alanya, Turkey, and forest management in Nepal [35][36]. Ostrom and her colleagues found that communities can and do create effective self-governance structures that prevent depletion, thereby debunking the idea that common resources are always over-exploited.
5.1   Digital Common Pool Resources
Ostrom’s ideas about polycentric governance can be adopted and adapted to digital assets and their governance. Here, we present the novel idea of “Digital Common Pool Resources” (DCPRs) as a digital adaptation of her theory of Common Pool Resources (CPRs). In this context, resources are not physical but digital, like codebases, online communities, or cryptographic assets. Ostrom’s principles find renewed relevance in the rapidly evolving world of blockchain and cryptocurrency, where decentralized, community-based control is paramount. Scholars such as De Filippi and Hassan [21] have further explored how blockchain technologies align with Ostrom’s principles, offering new opportunities for cooperative governance. Ostrom’s neoinstitutional approach provides a foundational framework for understanding self-governance in decentralized digital systems, focusing on the human-centered aspects of these technologies. This approach emphasizes collective action, reciprocity, and shared governance, aligning closely with the core philosophies of decentralization in the digital age.
To effectively translate Ostrom’s principles to digital assets, we must consider the nuances of decentralized governance. These principles, when reimagined for Web3, inform the design of resilient, equitable, and transparent systems. Among the challenges faced by digital commons is the cost and challenge of implementing security measures. More broadly, Ostrom’s principles provide guidance on creating sustainable digital communities:
  1. 1.
    Clearly defined boundaries: Digital assets should have clear ownership and access rights.
  2. 2.
    Proportional equivalence between benefits and costs: Participants should receive compensation proportional to their contributions.
  3. 3.
    Collective-choice arrangements: Community members should have a say in setting the rules.
  4. 4.
    Monitoring: The community should implement systems to monitor and validate transactions and interactions.
  5. 5.
    Graduated sanctions: Penalties for rule violations should vary based on the severity and context of the breach.
  6. 6.
    Conflict-resolution mechanisms: There should be accessible and fair processes for resolving disputes.
  7. 7.
    Minimal recognition of rights to organize: The community’s right to self-organize should be legally recognized.
  8. 8.
    Nested enterprises: Larger systems should be composed of smaller, interconnected subsystems with self-governance practices.
By applying these principles to the decentralized governance of Web3, the creation and management of digital commons can reflect the successful self-governance of tangible resources. Our approach, described below, helps mitigate the security challenges of digital spaces, ensuring that these commons are managed effectively and sustainably.
6   Graph Deep Learning to Identify Sybils
In decentralized governance systems, the presence of sybils poses substantial challenges for accurately measuring user influence, decision-making, and other aspects of community structure. These spurious identities can distort the representation of user roles, creating an inaccurate picture that also makes comparisons with other Online Communities literature problematic. This research seeks to address this challenge by developing a graph deep learning method tested against a popular DAO governance forum (snapshot.org). Specifically, the method involves a Graph Convolutional Neural Net (GCNN) with multiple engineered feature layers that learn vector node embeddings from a voting graph to construct a “similarity” subgraph, where similar nodes are identified using a fast vector search algorithm (Facebook AI Similarity Search, aka FAISS) and relabelled into clusters. The similarity graph is then reduced by combining predicted sybils.
6.1   Ethical considerations
It should be noted that using deep learning methods to analyze anonymous users on public financial and data records poses significant ethical risks and such methods should be considered a form of social network deanonymization. Even though the data are ostensibly public, individuals may not have consented or even be aware that their financial transactions could be linked back to them in an identifiable way. This becomes particularly troubling given that financial data can reveal sensitive aspects of an individual’s life—spending habits, income sources, and social networks—that they may not want publicly disclosed. In light of previous findings [15] on ethical research risks for cryptocurrencies, even when lacking real “human subjects,” research without consent exacerbates ethical concerns. Moreover, researchers undertaking this type of analysis must be acutely aware that their methods could be weaponized to infringe upon privacy or even perpetrate crimes.
6.2   Sybil Resistance and Defence in Depth for Community Design
The concept of anti-sybil mechanisms in community design is becoming increasingly vital. In a recent podcast [37], Gitcoin inventor Kevin Owocki and computer scientist Bryan Ford discuss efforts to create an identity system that counters the flaws of plutocratic governance, a particularly important issue given the widespread acceptance of coin voting in DAOs. Their recommended approach integrates principles from mechanism design and economic game theory, focusing on building identity systems that seek to preserve privacy yet provide durable digital personhood. As such, their proposed identity system would seek to provide strong anonymity and yet defend against sybils by leveraging social ties (trust networks), centralized authorities, or maybe even other cryptographically sophisticated mechanisms (such as Zero Knowledge Proofs).
These efforts, largely emerging from computer science, parallel the digital evolution of Ostrom-style commons management practices emerging in DAOs. DAOs continue to experiment with alternative identity systems, including Proof of Participation (PoP), Decentralized IDs (DIDs), and Self Sovereign IDs (SSIDs). New identity systems that guarantee digital personhood are promising but require complex, unproven infrastructure. This remains an active area of research and development.
The integration of blockchain technology in digital commons has brought about a paradigm shift in how commons security is addressed. Unlike traditional commons, which primarily focus on monitoring and surveillance to prevent illegal resource extraction, digital commons, particularly those managed by DAOs, face a different set of challenges. In response to these evolving security needs, our method does not necessitate structural changes to existing governance systems and requires no additional identity system. Instead, it emphasizes flexible, community-specified security responses to effectively counteract sybil attacks.
If sybils can be accurately and precisely identified, communities can collectively generate rules to deal with their behaviour. As such, effective sybil identification opens up a policy and design space. Ostrom, for instance, recommended leniency against community members who infringe the rules in traditional commons settings. Likewise, digital commons can be flexibly managed by identifying sybils and systematically modulating voting behaviours (by attaching negative karma scores, temporarily blocking spam voting, or even using a form of social naming and shaming). Importantly, our approach is permissive to current and de facto token voting schemes and could—if perfected—permit anonymous token voting while protecting against plutocratic outcomes.
However, this flexible security response raises a crucial question: what distinguishes a spam voter from a sybil voter, and is the distinction significant for DAOs? Conceptually, this distinction requires further analysis (is a spam voter just a sybil from the perspective of token voting?), but from a methodological standpoint, the distinction is pressing. If we define a sybil as a cluster of high dimensional embeddings, this implies a single hidden entity and a method for clustering based on similar vector embeddings seems appropriate. Conversely, if sybils are equated with spam (unwanted or harmful voting), targeting anomalous node embeddings for blocking becomes more relevant. Our approach adopts the former, differentiating between spam and sybils but more broadly it suggests a defense-in-depth strategy for DAOs, where outer layers target spam (anomalous voters) and inner layers focus on sybil threats (clusters of similar voters).
6.2.1 Defining Sybils and Deanonymizing Social Networks
There is a small but varied literature on sybils and social network deanonymization. The research often treats the problem of sybils and the deanonymization of social networks as separate, likely because scholars tend to shape their objectives around the methodologies they have at hand. For example, Douceur’s seminal work, “The Sybil Attack” [13], defines a sybil by its possible validation—indirectly, by either a trusted, centralised name authority or a web of trust—or directly, by an identity’s ability to control networked assets. This is a strict definition of sybils. For sybil defence in polycentric governance contexts, however, we can adopt a much less rigid definition. For us, sybils are clusters of distinctly similar anonymous voters.
We approach the challenge of identifying sybils as a graph signal problem where we exploit poor or low k-anonymity (wallet re-use) and use distinguishing metadata to construct higher dimensional representations to attack high k-anonymous (“singleton”) sybils. Sweeney’s [48] foundational theory of k-anonymity was initially designed to protect privacy by ensuring that any released data could not be traced back to less than k individuals. In the simplest terms, the theory is about blending in to stay anonymous. Some of our dataset is k=1 (singletons): 71% of unknown users reuse wallet addresses (see Figure 1), which we capture in our graph signal as multi-edge connectivity. These are the easier cases and despite a significant impact on privacy, most users on most chains reuse wallets (e.g., in Bitcoin, wallet reuse has hovered around 50-75%; (see Figure 2 from [6]). As explained below, our method to overcome k-anonymous datasets is deceptively simple: we generate node vector embeddings and cluster the high dimensional embeddings to identify similar, unknown voters. Importantly, since we generate (higher dimensional) node embeddings for all nodes with features, we bypass the protections of k-anonymity, making it possible to spot sybils more effectively.
Figure 1: Histogram of Wallet Reuse in Snapshot Voting Dataset. 71% of voters are $\geq$1.
Figure 2: Bitcoin Output Count Showing Wallet Reuse from BitMex Research [6].
Our method is far from novel, but it considers new contexts and uses sophisticated deanonymization techniques only recently made possible with advances in machine learning and the availability of powerful computational resources.
Prior research, like Backstrom, Dwork, and Kleinberg [4] tackled social network deanonymization, but (like Sweeney) used a dataset that was anonymized after-the-fact. The advantage of these approaches is that the method is supervised and has a clear objective. Unfortunately, such methods are not practical against real datasets that achieve their anonymity by cryptographically deriving new addresses. Their method fits their goal: they propose two graph-based techniques for deanonymization, an active one that requires the prior construction of a subgraph of attacking nodes (prior collusion) to reveal information about neighbours (by walking the graph or making strategic cuts), and a passive method that relies on graph isomorphisms. Narayanan and Shmatikov’s more recent work on social network deanonymization [34] draws inspiration from Backstrom et al., but used graph isomorphisms with auxiliary graphs to link anonymous identities with public identities, a more information rich context and a method that requires a small number of ‘seed’ nodes present in both graphs.
Narayanan and Shmatikov’s method for social network deanonymization is effective to the extent their walk-based propagation method can reach unknown nodes. However, since most anonymous networks (like Bitcoin) have many singleton nodes (assuming address rotation best security practices), propagation from known seeds to unknown nodes will not be effective.
Similarly, more recent work has been successful by using an open source information (OSI) approach to identify additional auxiliary graphs for deanonymizing Bitcoin users [7]. This research focuses on the ability to walk the transaction graph and uses a ‘follow the money’ approach to track down auxiliary information sources. However, in practical terms, this approach relies on poor security behaviour (high k-anonymity) and investigations often lead to an anonymizing coin mixer or a regulated exchange that can only be further investigated by legal authorities.
Recently, there have been significant attempts to deanonymize Bitcoin transactions. Most research uses supervised machine learning techniques to classify addresses, but with the exception of Akcora et al. [1] and Paquet-Clouston et al. [38], researchers do not have access to quality labelled datasets. Instead, researchers use synthetic and fake data: Ashfaq et al. [2] use a synthetic dataset; Rabieinejad et al. [40] generate fake labels; Dahiya et al. [10] use an unverified Kaggle dataset; Pham and Lee [39] use unverified labels for “30 thieves” of unknown provenance; Sankar Roy et al. [43] use the same Pham and Lee dataset. Or researchers use very simple heuristics (such as node degree patterns, see Weber et al. [51] and Lorenz et al. [31] who use the Weber et al. dataset) or slightly more complex heuristics (like motifs, see Wu et al. [54]) and assume that such patterns are evidence of complex criminal behaviour, like money laundering. This approach is a methodological dead end; putting aside the assumptions associated with fake, arbitrary, or false datasets, due to the imbalance of classes (very few labels of illegal or abnormal transactions), researchers must use resampling techniques like SMOTE to improve learning performance [54]. These are not good data science practices; effectively researchers are using low quality data and stretching it even thinner to boost learning performance (often using overly simple metrics like recall and accuracy). The results have no practical consequence. Indeed, surveying the field with all of these constraints and heuristics in mind, we can conclude that practical Bitcoin deanonymization remains elusive.
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [38]) or graph deep learning (Zhou et al. [57]), but these studies still have significant limitations. Paquet-Clouston et al. [38] attempts to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph, the latter uses the logic of sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use DBSCAN and k-means algorithms to cluster addresses and a naive similarity search. Finally, Zhou et al. [57] distinguish between address clustering and account identification, the former we call ‘sybil identification’ and the latter ‘deanonymization.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different context, is similar to the approach detailed below.
In the context of decentralized governance, however, sybils are identified using our more relaxed definition that circumvents k-anonymity. In this context, our method for identifying sybils is a weak form of deanonymization but one that reveals no personally identifiable information and unlike most previous methods does not require auxiliary or labelled graphs (although it can benefit from such information). Supervised techniques for deanonymization will not effectively identify sybils because, in a social graph characterized by strong anonymity, sybils may create a single, disconnected edge and thus lack sufficient connectivity to be identified by walk-based deanonymization approaches. Naive representations of sybils contain only 1 bit of structural information. Indeed, we assume that a majority of sybils are singletons. The existence of a labelled subgraph of known voters, linked to Ethereum Name Service (ENS) accounts, provides little help in identifying isolated sybils, for the same reason.
To overcome this limitation, we define a sybil as a cluster of “similar” nodes. Analytically, this makes sense; if nodes have similar higher dimensional representations and hidden features we can conclude that they are either representative of the same voting person or at least representative of voters acting in concert with one another (perhaps conducting a governance voting attack).
6.3   Experimental Sybil Identification Methodology
Our dataset covers governance voting for thousands of DAOs between 2020-07-18 and 2023-02-27 (959 days, 2.6 years), which covers approximately 7.2 million voting transactions. However, naive, exact counts of DAOs and their governance proposals are ineffective due to noise. For instance, upon manual inspection the dataset reveals many registered DAOs that failed to launch or were clearly set up as tests, and many governance proposals have durations far too short to be practical (see Figures 3 and 4).
Figure 3: Histogram of unfiltered governance proposal durations.
Figure 4: Histogram of filtered governance proposal durations.
A larger and more recent dataset is available, however due to the computational requirements of deep learning and the method’s broad effectiveness, we rarely analyzed the entire dataset. Instead, we found that fairly short timeframes (about one year) provided reasonable results and were computationally efficient.
We found evidence of extensive wallet address re-use, where a single voter makes multiple votes from the same pseudonymous wallet address (in the snapshot.org dataset, approximately 71% are reused wallets). In a traditional financial setting, this is considered poor security behaviour (resulting in low k-anonymity). However, in a polycentric governance context, this is not necessarily evidence of lax security practices but rather the demands of a different social context.
We performed typical Extract, Transform, and Load (ETL) practices on our dataset. After querying the snapshot.org API, we stored the dataset in a relational database and used a SQL command to order voting transactions by date. After sanitizing and checking the data, we constructed the voting graph. On the voting graph, we set node and edge attributes, including the amount of ‘voting power’ used for each vote on a proposal, UNIX timestamps, and persistent IDs linked to Ethereum Name Service IDs, when available. The resulting voting graph constructed from the snapshot.org dataset is available for inspection and statistical analysis (see Github). It is trivial to calculate statistical measurements like density and degree distribution, centrality measures like betweenness and eigenvector centrality, and community detection (there are no communities).
6.3.1 ‘Follow the money’ Sybil Chain Subgraph Depth-First Search
Using the strict definition of sybils [13], we can query and walk external graphs to search for sybil chains to support our self-supervised method. A sybil chain is a path across the graph that connects two or more unknown voters (singletons or k-anonymous nodes). To find such paths, we first construct multiple subgraphs. Specifically, we query the Covalent API for each wallet transaction on the origin token’s chain, construct subgraphs for each voter and connect these back to the original voter graph. Then, with this much larger graph we conduct a breadth-first search (BFS) for sybil chains. We use the cuGraph Python library for GPU BFS. We conclude that any path between two or more previously unknown voters is a sybil chain (a multiple-input clustering heuristic).
Figure 5: Examples of external subgraphs before attachment to voting graph.
Variants of this ‘follow the money’ approach are typical in the research, but as we discovered experimentally, without exceptionally large graphs the likelihood of finding sybil chains is very low (even in poor k-anonymous datasets, like Bitcoin, Ethereum, and our Snapshot). For testing purposes, we include a random fake sybil in the dataset and identify it.
Figure 6: Small voting graph with fake sybil chain identified using BFS.
6.3.2 Graph Convolutional Neural Network (GCNN)
With the voting graph, we engineer learnable features and train a Graph Convolutional Neural Network (GCNN) to create vector embeddings for each node. Importantly, this method is not limited to extensive wallet re-use (which we do see), rather, because the method generates embeddings for all nodes, including singletons, it is potentially effective against all graphs, including k-anonymous ones.
There are any number of possible custom Graph Convolutional Neural Networks (GCNNs) tailored for node classification tasks on graphs, with a multitude of interchangeable layers for various tasks. For our method, the first few steps are relatively straightforward. First, construct a voting graph where known (labelled) voters and re-used wallets are combined into single nodes with multiple edges to proposal nodes (we use a multi- or hyper-graph to accommodate multiple edges from each node). Second, we engineer features to learn higher dimensional vector embeddings on all nodes. We use a simple reconstruction loss (Mean Squared Error or MSE) for training the GCNN. Third, we use the node embeddings to conduct a k-means vector clustering on unknown nodes. We then propagate the most common label associated with each cluster to relabel the actual, lower dimensional nodes to form the sybil cluster. Finally, we take the relabelled similarity graph and combine nodes with their labels in each cluster to form a higher dimensional, reduced-size graph. The resulting graph is an inductive prediction of sybil nodes.
The deep learning method was developed with the Python libraries NetworkX and RAPIDS AI to construct the graphs, the PyTorch Geometric library for the GCNN, and the Facebook AI Similarity Search (FAISS) for the fast embeddings clustering. Whenever possible, GPU accelerated libraries were used, although speed and memory tradeoffs quickly become a practical data analysis challenge when dealing with large graphs requiring several hundred gigabytes of memory for large matrix multiplications during learning. As an inductive method, decisions to use a particular library, feature, or machine learning algorithm were made pragmatically and tested against the dataset.
We develop features and construct a typical train/test/validation split for the forward pass of our GCNN. The first layer of the forward pass is fully connected to capture node features and to incorporate voting edges and their ‘power’ attributes; the second layer incorporates the edge and node features with a sequential multi-layer perceptron (MLP) with ReLu activation that is then passed into a convolutional layer with mean feature aggregation; the third layer utilizes temporal sequences that have been engineered for a Long Short Term Memory (LSTM) layer; and finally, the fourth layer uses a multi-headed Graph Attention Network (GAT) to search and better learn edge connectivity. Each of these layers has tuneable hyperparameters that were set by conducting a grid search. After each forward pass and backpropagation to set training weights, training loss (MSE) was calculated, visualized, and assessed. The resulting tensors are normalized on zero and were checked for sufficient numeric variability.
The four layer GCNN to identify sybils:
  1. 1.
    Fully Connected (FC) Layer with voting power
  2. 2.
    Multi-Layer Perceptron (MLP) Convolutional Layer with node features
  3. 3.
    Long Short Term Memory (LSTM) with temporal features
  4. 4.
    Multi-Headed Graph Attention Network (GAT) with edge index.
6.3.3 Fast Vector Embeddings Search using FAISS
Once we have learned higher dimensional voter and voting representations for all (known and unknown) nodes, we can cluster similar vector embeddings, which we initially learned using the GCNN. In deciding on our method, we also considered other similarity metrics like Euclidian distance, cosine similarity, and Jaccard similarity, but computing many-to-many similarity searches on large matrixes is very slow $O(n^{2}\cdot d)$. Thankfully, the Facebook AI Similarity Search (FAISS) Python library provides extremely fast and easily scaled search and clustering options (including GPU processing). FAISS offers several search index options for improving efficiency, including approximate search methods that can scale to billions of elements. For our purposes, the L2 index was quick enough and provides exact results (for much larger datasets, FAISS provides approximate methods that are substantially quicker). We use the included k-means clustering method to find sybil clusters, which uses centroids in the high dimensional embeddings to cluster and return neighbours. We normalize the clusters by dropping any remaining singletons and large clusters (mean + 1 standard deviation).
Figure 7: Normalized Sybil Cluster Sizes.
We then inspect each cluster and propagate the most common label across the cluster to construct a “similarity” graph. Finally, we construct a reduced “clustered” graph by combing clustered nodes (and reconstructing edges).
Description Value
Total clusters formed (excluding singletons) 655
Average cluster size 131.95 nodes
Largest cluster contains 41,149 nodes
Smallest cluster contains 2 nodes
Filtered clusters formed (excluding large clusters) 651
Largest cluster after filtering contains 1,877 nodes
Table 1: Summary of Cluster Analysis for 300,000 Votes
In formal notation:
  • $X$: Node features matrix (degree features).
  • $E$: Edge features matrix (voting power).
  • $T$: Temporal features matrix.
  • $A$: Adjacency matrix of the graph.
  • $W_{\text{FC}},b_{\text{FC}}$: Weights and bias of the Fully Connected layer.
  • $W_{\text{NN}},b_{\text{NN}}$: Weights and bias of the NNConv layer.
  • $W_{\text{LSTM}},h$: Weights and hidden state of the LSTM layer.
  • $W_{\text{GAT}}$: Weights of the GAT layer.
  • $D$: Embedding dimension for FAISS L2 index.
The GCN model can be expressed as:
[…]
$\quad\text{ReLU}(XW_{\text{FC}}+b_{\text{FC}}),$
$\quad A,E;W_{\text{NN}},b_{\text{NN}}\Big{)},$
$\quad T;W_{\text{LSTM}},h\Big{)},A;W_{\text{GAT}}\Big{)}$
The output of this GCN model is a set of node embeddings, denoted as Emb, where
Emb $=\text{GCN}(X,A,E,T).$
The FAISS L2 index is then used to efficiently search within these embeddings. Let Query be a query vector of dimension $D$. The FAISS L2 index searches for the nearest neighbors of Query in the embedding space:
[…]
$\text{FAISS}_{L2}(\text{Query},\text{Emb},D)$
7   Results
We have conducted numerous graph analysis experiments on DAO governance voting with a range of graph sizes, learnable features, specialized convolutional layers, and training and optimization algorithms. We also tested different clustering approaches including DBMScan, spectral clustering, and unsupervised methods; we deployed supervised classification methods with the labelled subgraph; and we considered the problem from different graph analysis viewpoints (such as considering sybil identification as a problem of link prediction using graph autoencoders, and as a community detection problem using Girvan-Newman and Louvain). While many of these methods initially looked promising, they ultimately failed for a variety of reasons.
The most common reason for experimental failure was the lack of edge connectivity on sybils, which is precisely what provides anonymity. However, our task of identifying sybils is easier than complete deanonymization and by approaching the problem as a k-anonymous graph signal we might achieve what we want anyway, which is the ability to defend against attackers at scale but not reveal individual identities. This is an easier, but not easy problem. Since we should expect sybils to be singletons, also like voting spam, we cannot effectively cluster their lower dimensional structure, nor can we use a labelled subgraph of known users for supervised learning, and nor will we be able to detect communities or cliques. Instead, we defined sybils as clusters of similar higher dimensional vector node embeddings. And since we form clusters using a k-means clustering method on higher dimensional embeddings, singletons pose no particular challenge for our method.
Tested against the snapshot.org dataset, our method typically reduces the original voting graph by approximately 2-5% (subject to hyperparameters). The method produces a voting graph with real IDs that can be inspected for further forensic analysis or visualized (see Figures 8 and 9). Moreover, we can produce a few key sociometric measurements to begin the process of constructing community baselines for further research (see Table 2).
Metric Value
Start Date 2020-07-18
End Date 2021-04-20
Duration 0 years and 275 days
Original Graph 89,833 nodes, 300,000 edges
Similarity Graph 89,833 nodes, 300,000 edges
Clustered Graph 87,874 nodes, 300,000 edges
Number of Known Voters 11,510
Number of Unknown Voters 69,150
Number of Potential Sybils Identified 9,947
Number of Sybil Clusters Formed 301 (0.35% of Unknown Voters)
Node Reduction After Clustering Sybils 3.7%
Table 2: Sociometric Results for 300,000 Votes
Figure 8: 9,497 Potential Sybils Identified in 651 Clusters for 300,000 Votes.
Figure 9: 301 Sybils Combined for a 3.7% Reduction in Voters.
8   Future Research
The emergence of Decentralized Autonomous Organizations (DAOs) marks a pivotal shift in the governance of digital assets. DAOs, operating without centralized control, face significant challenges from sybil attacks, where one entity creates many identities to gain disproportionate influence. Building on Ostrom’s principles for managing commons, we recognize that leniency might not directly apply in the digital realm due to the anonymous nature of offenders. Consequently, once armed with effective tools to identify sybils, DAOs would still need to experimentally discover governance protocols to mitigate sybil threats at scale. Penalties must be carefully calibrated to deter malicious behaviour without stifling the collaborative ethos that underpins DAOs.
To improve and further validate our inductive method of sybil identification we need benchmarks from other online communities. Moreover, more sophisticated and further development of machine learning features would enhance the method. Finally, much more computationally expensive ‘following the money’ analyses that trace coins further to their origins would add a valuable graph signal for supervised learning.
In addition to refining current methods, it is worth exploring other deep learning techniques. To adapt the method for detecting voting spam, we could instead detect anomalies within the node embeddings. Thus, instead of identifying similar nodes, this approach searches for outliers that could indicate spam activity. In coin voting contexts, where the integrity of each vote is tied to a digital token, distinguishing between voting anomalies (spam) and similarities (sybils) is critical for safeguarding the democratic processes within DAOs.
9   Conclusion
The Internet’s transformative impact on communication costs has laid the groundwork for a similar paradigm shift in governance technologies. By drastically reducing the costs associated with dispute resolution and collaboration, these technologies have the potential to alter the landscape of governance. This shift is frustrated by anonymous accounts called sybils, but by deploying a deep learning method we augmented network transparency to minimize the costs of monitoring and enforcement within governance structures. Without impacting personal privacy and without relying on unproven identity systems, our approach modulates governance behaviours at scale to prevent sybil attacks, potentially heralding a new era of efficiency and effectiveness in governance systems.
The concept of Digital Common Pool Resources (DCPRs) and polycentric governance emerges as a promising avenue for future research in this context. This governance model is particularly relevant in addressing the perennial challenges associated with managing digital assets in a commons, such as congestion, overuse, pollution, and destruction. Traditional approaches often struggle with enforcing limits on resource utilization, but polycentric governance offers a novel framework that could be more effective in managing these resources. With polycentric governance, participants in the digital commons develop context-specific rules to prevent infringement and establish consequences for those who exploit the commons. By leveraging decentralized decision-making and consensus mechanisms, these regimes hold the potential for more sustainable and equitable resource management.
The emerging field of DAO Studies provides insightful parallels for polycentric governance. DAOs operate on principles of distributed authority and decision-making, which align closely with the tenets of polycentric governance. This alignment offers a unique opportunity to draw insights from DAOs to inform traditional policy, regulation, and governmental contexts. The lessons learned from DAO governance models, such as consensus building, decentralized control, and community-driven decision-making, could offer valuable strategies for tackling governance challenges in a broader context. By integrating these insights, traditional governance systems could evolve to become more adaptive, participatory, and responsive to the needs of their constituencies.